# Effective Hamiltonian for Twisted TMDs which Exhibit Pressure Dependent Topological Phase Transitions

### v1

@@ III.1 Untwisted case
One can model this kind of pressure in DFT using the Abinit software [62] by setting the value of the $z$-component of the target stress tensor $\tau_{zz}$ in the structural optimization computation. We note that pressure from the stress tensor can then be calculated as
@@
To later obtain pressure-dependent coupling constants for our effective Hamiltonian, band structures for a range of pressures were needed. For each band structure we computed, we relaxed both the unshifted and shifted structures under pressures ranging from 0 - 4 GPa. Our relaxations were based on the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [63, 64, 65, 66]. The computations were performed under the fully relativistic scheme using PAW pseudopotentials [67] and Perdew-Wang LDA exchange-correlation functional [68, 69]. The wave functions energy cutoff was set to 24 Ry, which ensured that the self-consistent field (SCF) step converged. Moreover, we chose a $\Gamma$-centered $16\times 16\times 1$ grid of $k$-points for the SCF calculations. The tolerance on the potential residual was set to $2\times 10^{-12}$ Ry as the convergence criterion for the SCF calculation, whereas the convergence for the relaxation was based on the maximum absolute force on the atoms, which was set to $2.0\times 10^{-4}$ Ry/Å.
@@
respectively. We also obtain the inter-layer distance of $d_{0}$ and $d_{1}$, $D_{z}^{(d_{0})}$ and $D_{z}^{(d_{1})}$, as
@@
The DFT band structures of both $d_{0}$ and $d_{1}$ systems at various pressures are shown in Fig. 10.
A vital idea we checked in our computations was that the honeycomb structure of the lattice for both shifted systems was kept intact during the relaxation (at least for our pressure range). Moreover, we made sure that the two topmost valence bands remained quadratic curves near the $\mathbf{K}$ valley (this was true up to 3.5 GPa). This property enabled us to adapt our low-energy Hamiltonian Equation (21) to structures under pressure. We note that as seen in Fig. 10, the quadratic behavior of the topmost valance bands breaks down at the pressure of 3.6 GPa for the $d_{0}$ structure and 3.9 GPa for the $d_{1}$ structure. Based on this result, we stress that our model is only reliable for pressure in the 0.0 - 3.5 GPa range.
Figure 11: (a, c, d, e) Plots of the coupling parameters as function of pressure. The data is marked as red dots, while the blue line is our fit function. (b) The energy difference between the topmost valence bands of $d_{0}$ and $d_{1}$ structures.
As in Sec. II.1, we can fit the energy data of the region around the $\mathbf{K}$ valley obtained from DFT computations to the effective Hamiltonian (21) to obtain the coupling parameters at each pressure. The plot of coupling $V$ as a function of pressure is shown as red dots in Fig. 11(a). The data could be fitted very well to a function of the form:
@@
III.2 Twisted Case
@@
III.2.1 Twist Angle $\theta=1.4^{\circ}$
@@
Fig. 13(e) shows the bandwidth of the three topmost bands. As can be seen, pressure affects the bandwidth of those bands. As an exciting sample, the width of the first band $\epsilon_{w,1}$ is less than 1 meV for pressure ranging from 0 – 1.4 GPa. Although it increases to the maximum of 1.8 meV within the range of 1.4 – 2 GPa, the bandwidth becomes narrow again after 2 GPa and reaches its minimum width of 0.12 meV at pressure $P=3.5$ GPa as can be seen in Fig. 12. This type of band (with bandwidth $<1$ meV), usually called an ultra-flat band, might indicate a strong correlation for electrons occupying the corresponding band.
III.2.2 Twist Angle $\theta=2.3^{\circ}$
@@
Next, we investigate the case of twist angle $\theta=2.3^{\circ}$ which in the unpressurized case is the angle where the topological transition happens. As shown in 14(a-c), there are several transitions for the second and third bands. The first transition occurs immediately at $P=0.1$ GPa after both bands touch at the $\bm{\gamma}$ point where the valley Chern numbers of the second and third bands change from $(C_{2},C_{3})=(-1,2)$ to $(C_{2},C_{3})=(1,0)$. Both bands cross again at $P\sim 1.3$ GPa, as shown in Fig. 14(d), and the band crossing is accompanied by a change in valley Chern numbers from $(C_{2},C_{3})=(1,0)$ to $(C_{2},C_{3})=(-1,2)$. Another band crossing occurs at the same point in $k$ space, i.e., $\gamma$, at pressure $P\sim 3.1$ GPa. Here, the band crossing is followed by a change of valley Chern numbers as $(C_{2},C_{3})=(-1,2)\to(1,0)$. Although both bands seem to cross in Fig. 15, the gap between the second and third bands is 0.03 meV. A final change of valley Chern numbers occurs after the crossing at $P\sim 3.3$ GPa as $(C_{2},C_{3})=(1,0)\to(-1,2)$.
@@ IV Conclusion
In summary, we started our work with a demonstration of the procedure to obtain the effective low-energy Hamiltonian for the twisted TMD based on a general symmetry-based approach [54] and reproduced results that closely resemble those of Ref. [22]. We then use the energy data for pressurized untwisted TMDs from DFT computation to obtain an explicit form of the physical coupling parameters $V,\psi,w$ and $m^{*}$ as functions of pressure $P$ which are summarized in Eq. (33)–(36). Using these parameters, we have also calculated the valley Chern numbers of the topmost three bands and observed multiple topological transitions as demonstrated by the changes in valley Chern number for the second and third topmost bands. These changes occur at band crossings when the pressure reaches certain critical values. The occurrence of these transitions means that varying pressure can have a similar effect as varying the twist angle in the sense that both procedures give rise to non-trivial valley Chern numbers. Variations in pressures could be advantageous compared to changes in twist angle since an experimentalist could modify the material properties in situ without the need to prepare different materials for each different twist angle to obtain the non-trivial topological insulating property.
@@ V Acknowledgements
M. V. and H. B. gratefully acknowledge the support provided by the Deanship of Research Oversight and Coordination (DROC) at King Fahd University of Petroleum & Minerals (KFUPM) for funding this work through exploratory research grant No. ER221002.

### v4

@@ III.1 Untwisted case
One can model this kind of pressure in DFT using the Abinit software [67] by setting the value of the $z$-component of the target stress tensor $\tau_{zz}$ in the structural optimization computation - other components have their target set to zero. We note that pressure from the stress tensor can then be calculated as
@@
To obtain pressure-dependent coupling constants for our effective Hamiltonian, band structures for a range of pressures were needed. For each band structure we computed, we relaxed both the unshifted and shifted structures under pressures ranging from 0 - 4 GPa. Our relaxations were based on the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [68, 69, 70, 71]. The computations were performed under the fully relativistic scheme using PAW pseudopotentials [72] and Perdew-Wang LDA exchange-correlation functional [73, 74]. The wave functions energy cutoff was set to 24 Ry, which ensured that the self-consistent field (SCF) step converged. Moreover, we chose a $\Gamma$-centered $16\times 16\times 1$ grid of $k$-points for the SCF calculations. The tolerance on the potential residual was set to $2\times 10^{-12}$ Ry as the convergence criterion for the SCF calculation, whereas the convergence for the relaxation was based on the maximum absolute force on the atoms, which was set to $2.0\times 10^{-4}$ Ry/Å.
@@
respectively. We notice that although both $a_{0}$ are different, the difference is insignificant (less than 0.4% for the pressure range of interest). Thus, for the calculation in section III.3, we will approximate using $a_{0}^{(d_{1})}(p)$. This also agrees with previous work from the literature in the unpressurized case that also neglected these differences. This step, we stress, is important because it considerably simplifies our analysis.
We also obtain the inter-layer distance of $d_{0}$ and $d_{1}$, $D_{z}^{(d_{0})}$ and $D_{z}^{(d_{1})}$, as
@@
With the relaxed structural parameters, we then compute the band structure in Quantum Espresso using the same simulation parameters as discussed in Sec. II.1. The DFT band structures of both $d_{0}$ and $d_{1}$ systems at various pressures are shown in Fig. 10.
A vital idea we checked in our computations was that the honeycomb structure of the lattice for both shifted systems was kept intact during the relaxation (at least for our pressure range). Moreover, we made sure that the two topmost valence bands remained quadratic curves near the $\mathbf{K}$ valley (this was true up to 3.5 GPa). This property enabled us to adapt our low-energy Hamiltonian Equation (21) to structures under pressure. We note that as seen in Fig. 10, the quadratic behavior of the topmost valance bands breaks down at the pressure of 3.6 GPa for the $d_{0}$ structure and 3.9 GPa for the $d_{1}$ structure. Based on this result, we stress that our model is only reliable for pressure in the range of 0.0 - 3.5 GPa.
Another important observation is that after a certain critical pressure, bands in the $\bm{\Gamma}$ region enter our energy window of interest. A more detailed analysis of the $\bm{\Gamma}$ region, would therefore be needed to capture the complete physical picture at such pressure. This observation offers a great opportunity for a follow-up work with exciting physics - for instance, there could be an exciting interplay between $\bm{\Gamma}$ and $\mathbf{K}$ valley electrons that is mediated by short-range deformations. However, this is beyond the scope of the current work, where we mostly focus on electronic contributions from the $\mathbf{K}$ point region.
Figure 11: (a, c, d, e) Plots of the coupling parameters as function of pressure. The data is marked as red dots, while the blue line is our fit function. (b) The energy difference between the two topmost valence bands of $d_{0}$ and $d_{1}$ structures. Since parameter $V$ corresponds to the gap between the two bands, its oscillation-like behavior originates from the oscillation of the energy gap with pressure variations.
As in Sec. II.1, we can fit the energy data of the region around the $\mathbf{K}$ valley obtained from the DFT computations to the effective Hamiltonian (21) to obtain the coupling parameters at each pressure. The plot of the coupling parameter $V$ as a function of pressure is shown as red dots in Fig. 11(a). The data could be fitted very well to a function of the form:
@@
We emphasize that the pressure-dependent couplings given above constitute one of the key results of our paper. Moreover, it is important to note that while the result in the current paper is primarily used to study twisted systems, its application extends beyond this. Indeed, the approach can be used, for instance, to model position-dependent pressure configurations—such as those obtained by pressing on the system via a needle. Similarly, one could use the approach to study bilayers with partial dislocations and pressure acting simultaneously. Of course, it is important to ensure that deformations happen on large enough distance scales such that they do not allow electrons to tunnel from the K valley to other valleys. The model, therefore, should be useful in an understanding of mesoscopic physics.
III.2 Contribution from Beyond K Region
As already shown in Figure 10, bands near the $\bm{\Gamma}$ point enter the energy range of interest at higher pressure. To better understand contributions to the physics from the $\mathbf{K}$ valley region we analyze the density of states (DOS) of the system.
Here, after performing an SCF calculation using Quantum Espresso like in Section III, we compute the energy of the $d_{0}$ structure at different pressures. As representative examples, we discuss the cases for pressures 0.0 and 1.8 GPa. Due to a sixfold rotational symmetry, we can focus only on one segment of the Brillouin zone, for example, the region enclosed by $\mathbf{K}-\bm{\Gamma}-\mathbf{K^{\prime}}$ as seen in Fig. 12. We compute the energy for 5151 uniformly distributed $k$-points of this region and then divide the energy data into three subsets; the energy that came from the $\mathbf{K}$ region, the $\bm{\Gamma}$ region, and outside of both regions. We include all points within the radius of 0.1 (in the unit of $2\pi/a_{0}$) from $\mathbf{K}$ and $\bm{\Gamma}$.
Figure 12: The region used for computation of the density of states. We take 5151 uniformly distributed points in this region for our DOS calculation. Then, we divide them into three subsets: $\mathbf{K}$ (and $\mathbf{K^{\prime}}$) region, $\bm{\Gamma}$ region, and the outside region.
The density of states at energy $E$ is then defined as
$g(E)=\sum_{n,\mathbf{k}}\delta(E-E_{n}(\mathbf{k})),$ (37)
where $\delta$ is the Dirac delta function, and $E_{n}(\mathbf{k})$ is the energy of the $n$-th state at momentum $\mathbf{k}$.
Figure 13: The density of states plot of $d_{0}$ structure at 0.0 and 1.8 GPa with the region marked in blue is the low-energy region where the quadratic momentum terms in our effective Hamiltonian provide a good approximation. At 0.0 GPa, the $\mathbf{K}$ has the dominant contribution while at 1.8 GPa it gives about 29.9% contribution to the total DOS. Also, for both cases, the contribution from the $\bm{\Gamma}$ region is equal to zero in this energy region.
To obtain an interpretable plot, it is customary to approximate the Dirac delta function with a Gaussian function,
$\delta(E-E_{n}(\mathbf{k}))\approx\frac{1}{\sqrt{2\pi}\sigma}\exp{\left(-\frac{(E-E_{n}(\mathbf{k}))^{2}}{2\sigma^{2}}\right)}$ (38)
where $\sigma$ is a smearing parameter. The total DOS of the system and the contributions from each region are shown in Figure 13.
The plot of the DOS reveals that at zero pressure, the dominant contribution for the DOS in the energy range that we are interested in comes from the region around $\mathbf{K}$. This means that many physical properties of the system can be expected to have a large contribution from electrons with momentum in this region of the Brillouin zone. However, at higher pressure, the contribution from this region becomes less dominant - although it remains important which amounts to around 29.9%. We also note that for both cases, the contribution from the $\Gamma$ region is virtually zero in the energy range of our interest. The result for the $d_{1}$ structure also shows a similar trend. Hence, further analysis of the contribution from other regions is needed to capture the complete physics of the system. Although our simplified model could not capture the full story of the system, it is expected to provide significant insights into parts of the physics - most importantly $\mathbf{K}$ valley Chern numbers, which can be useful in the valleytronics context [75, 76].
III.3 Twisted Case
@@
III.3.1 Twist Angle $\theta=1.4^{\circ}$
@@
Fig. 15(e) shows the bandwidth of the three topmost bands. As can be seen, pressure affects the bandwidth of those bands. As an exciting sample, the width of the first band $\epsilon_{w,1}$ is less than 1 meV for pressure ranging from 0 – 1.4 GPa. Although it increases to the maximum of 1.8 meV within the range of 1.4 – 2 GPa, the bandwidth becomes narrow again after 2 GPa and reaches its minimum width of 0.12 meV at pressure $P=3.5$ GPa as can be seen in Fig. 14. This type of band (with bandwidth $<1$ meV) can be considered an ultra-flat band[77]. Flat bands are interesting because they might indicate the relevancy of strong correlation effects for electrons occupying the corresponding band. The literature discusses various mechanisms that can lead to flat bands in twisted moire systems[78, 49, 79]. Since the Hamiltonian discussed in [49] is a close match to our Hamiltonian, one can expect that their description of a flat-band mechanism also applies to our work.
III.3.2 Twist Angle $\theta=2.3^{\circ}$
@@
Next, we investigate the case of twist angle $\theta=2.3^{\circ}$ which in the unpressurized case is the angle where the topological transition happens. As shown in 16(a-c), there are several transitions for the second and third bands. The first transition occurs immediately at $P=0.1$ GPa after both bands touch at the $\bm{\gamma}$ point where the valley Chern numbers of the second and third bands change from $(C_{2},C_{3})=(-1,2)$ to $(C_{2},C_{3})=(1,0)$. Both bands cross again at $P\sim 1.3$ GPa, as shown in Fig. 16(d), and the band crossing is accompanied by a change in valley Chern numbers from $(C_{2},C_{3})=(1,0)$ to $(C_{2},C_{3})=(-1,2)$. Another band crossing occurs at the same point in $k$ space, i.e., $\gamma$, at pressure $P\sim 3.1$ GPa. Here, the band crossing is followed by a change of valley Chern numbers as $(C_{2},C_{3})=(-1,2)\to(1,0)$. Although both bands seem to cross in Fig. 17, the gap between the second and third bands is 0.03 meV. A final change of valley Chern numbers occurs after the crossing at $P\sim 3.3$ GPa as $(C_{2},C_{3})=(1,0)\to(-1,2)$.
We expect that these pressure-induced changes in Chern number would manifest experimentally as pressure-induced changes in the Hall current for materials with appropriate band filling.
@@ IV Conclusion
In summary, we started our work with a demonstration of the procedure to obtain the effective low-energy Hamiltonian for the twisted TMD based on a general symmetry-based approach [55] and reproduced results that closely resemble those of Ref. [22]. We then use the energy data for pressurized untwisted TMDs from DFT computation to obtain an explicit form of the physical coupling parameters $V,\psi,w$ and $m^{*}$ as functions of pressure $P$ which are summarized in Eq. (33)–(36). The model that we developed here is valid around the $\mathbf{K}$ region and can capture some physics of the system under pressure and further development is needed to obtain the contribution from other regions.
Using these parameters, we have also calculated the valley Chern numbers of the topmost three bands and observed multiple topological transitions as demonstrated by changes in valley Chern numbers for the second and third topmost bands. These changes occur at band crossings when the pressure reaches certain critical values. The occurrence of these transitions means that varying pressure can have a similar effect as varying the twist angle in the sense that both procedures give rise to non-trivial valley Chern numbers. Variations in pressures could be advantageous compared to changes in twist angle since an experimentalist could modify the material properties in situ without the need to prepare different materials for each different twist angle to obtain the non-trivial topological insulating property
@@ V Acknowledgements
M. V. and H. B. gratefully acknowledge the support provided by the Deanship of Research Oversight and Coordination (DROC) at King Fahd University of Petroleum & Minerals (KFUPM) for funding this work through exploratory research grant No. ER221002. The Alfahd High-Performance Computing facility, provided by the College of Petroleum Engineering & Geosciences at KFUPM, is acknowledged for its essential role in supporting the computational calculations.
VI Data Availability
Data that supports the findings of this article is available from the authors upon reasonable request.